### v3

# Aligning NLP Models with Target Population Perspectives using PAIR: Population-Aligned Instance Replication

Abstract
Models trained on crowdsourced annotations may not reflect population views, if those who work as annotators do not represent the broader population. In this paper, we propose PAIR: Population-Aligned Instance Replication, a post-processing method that adjusts training data to better reflect target population characteristics without collecting additional annotations. Using simulation studies on offensive language and hate speech detection with varying annotator compositions, we show that non-representative pools degrade model calibration while leaving accuracy largely unchanged. PAIR corrects these calibration problems by replicating annotations from underrepresented annotator groups to match population proportions. We conclude with recommendations for improving the representativity of training data and model performance. [^1] The code for experiments is available at https://github.com/soda-lmu/PAIR.
Aligning NLP Models with Target Population Perspectives using PAIR:
Population-Aligned Instance Replication
[…]
1 Introduction and Inspiration
When a hate speech detection model flags harmless expressions as toxic, or a content moderation system fails to identify genuinely harmful content, the root cause often lies not in the model architecture, but in who annotated the training data. While Natural Language Processing (NLP) models aim to serve broad populations, the human judgments used to train these systems often come from crowdworkers and convenience samples. And the demographics, cultural contexts, and worldviews of these annotators often differ from those of the communities the models ultimately impact Sorensen et al. (2024); Fleisig et al. (2024). These non-representative annotator pools can have real consequences, because annotator characteristics like age, education level, and cultural background impact how content is annotated Sap et al. (2022); Fleisig et al. (2023); Kirk et al. (2024). The influence of annotator characteristics underscores that language understanding is not a single objective truth but a constellation of equally valid interpretations anchored in different lived experiences. When this perspectivist interpretation is ignored, models trained on non-representative data can perpetuate the biases and blind spots of their limited training data Berinsky et al. (2012); Hebert-Johnson et al. (2018); Mehrabi et al. (2021); Rolf et al. (2021); Hüllermeier and Waegeman (2021); Ouyang et al. (2022); Favier et al. (2023); Smart et al. (2024).
Figure 1: Top: Adjusting survey data to match population produces high quality results.
Bottom: Can a similar adjustment in data annotations also improve model performance?
Fortunately, survey researchers have developed statistical techniques to produce population-level estimates from non-representative samples Bethlehem et al. (2011). The top panel of Figure 1 shows a simple survey workflow: collecting survey data, creating statistical weights to match the sample to the population, and estimating population parameters. We adapt this approach to the machine learning context, enabling models to better align with target populations even when trained on non-representative annotator pools (bottom panel).
Our Population-Aligned Instance Replication (PAIR) method post-processes training data to better reflect target populations without collecting additional annotations. We test the approach with a simulation study (Burton et al., 2006; Valliant, 2019; Morris et al., 2019) and answer two questions:
  • RQ1: How do non-representative annotator pools impact model calibration and accuracy?
  • RQ2: Can our proposed weighting method (PAIR) mitigate these annotator pool effects?
Our results demonstrate that models trained on non-representative annotator pools perform worse than those trained on representative pools. However, simple adjustment methods can improve performance without collecting additional data. These findings suggest that insights from survey methodology can help artificial intelligence (AI) systems better represent the populations they serve.
2 Related Work
Several strands of related work inform our approach to identifying and mitigating bias due to the use of non-representative annotators:
Annotator Impact on Data and Models.
Annotator characteristics and attitudes significantly influence annotation quality, particularly for subjective tasks like toxicity detection (Giorgi et al., 2025; Prabhakaran et al., 2021; Fleisig et al., 2023; Sap et al., 2022). For example, annotators’ political views and racial attitudes affect their toxicity judgments (Sap et al., 2022). Models trained on non-representative annotator pools inherit these biases and generalize poorly (Berinsky et al., 2012; Mehrabi et al., 2021; Rolf et al., 2021; Ouyang et al., 2022; Favier et al., 2023; Smart et al., 2024; Mokhberian et al., 2024).
Annotator Demographics.
Several researchers advocate collecting annotator demographics to assess representation and identify biases (Bender and Friedman, 2018; Prabhakaran et al., 2021; Plank, 2022; Wan et al., 2023; Santy et al., 2023; Pei and Jurgens, 2023). [^2] In our context, these characteristics are used only to analyze bias. Because they are not available for unannotated text, they are not features that the model can use. However, collecting and releasing these data can raise privacy concerns (Fleisig et al., 2023). Recent works have also used demographics to prompt the large language models Argyle et al. (2023), and some find that these are less effective in subjective contexts Sun et al. (2025); Orlikowski et al. (2025).
Debiasing & Data Augmentation Methods.
Prior work has proposed various approaches to reduce bias in training data features and annotations. Most similar to our work is the resampling and reweighting approaches of Calders et al. (2009) and Kamiran and Calders (2012), imputation Lowmanstone et al. (2023), and the oversampling of minority class cases of Ling and Li (1998). PAIR adapts these methods to balance annotator characteristics rather than class labels or sensitive observation-level features. PAIR retains the simplicity and interpretability of earlier resampling methods while extending them to a “Learning with Disagreement” (Uma et al., 2021; Leonardelli et al., 2023) setting with multiple annotations per observation, by replicating annotations from underrepresented annotator groups.
3 PAIR Algorithm: Adjustment via Pseudo-Population
To adjust an annotator pool to better reflect a target population, we propose the PAIR algorithm, which constructs a pseudo-population through post-stratification, weight normalization, and deterministic replication. This adjustment strategy is inspired by established methods in survey sampling (Quatember, 2015).
Post-stratification aligns a sample more closely with population-level distributions (Bethlehem et al., 2011; Valliant et al., 2013). Annotators are grouped into strata based on demographic or behavioral characteristics. For each unit $i$ in stratum $s$, a post-stratification weight is computed as:
$w_{s,i}=\frac{P_{s}}{S_{s}}$ (1)
where $P_{s}$ and $S_{s}$ denote the share of the population in stratum $s$ and the share of the sample (or annotator pool), respectively. The $P$ values come from official statistics or surveys. The $S$ values likely come from the annotators themselves and researchers may have to collect them. This technique can accommodate multiple stratification variables; it is only limited by the availability of population or reference data and data about the annotators.
These weights have only relative meaning and are invariant to multiplication by a constant ($K$):
$w_{i}^{\text{normalized}}=w_{i}^{\text{initial}}\times K$ (2)
Normalization useful if research teams have a target number of annotations per observation in mind, for either computational or design reasons, or if some weights given by Eq. 1 are very small and round to one.
To generate a pseudo-population, we apply deterministic replication: each unit is replicated $n_{i}$ times where
$n_{i}=\text{round}(w_{i}^{\text{normalized}})-1$ (3)
ensuring integer replication counts. This approach produces a dataset that reflects population proportions while maintaining interpretability and reproducibility.
While we focus in this initial study on deterministic replication, alternative implementations are possible, including resampling-based replication or direct incorporation of weights into model training.
4 Annotation Simulation and Model Training
To address our research questions, we conduct a simulation study on offensive language and hate speech detection. We imagine a population made up of equal shares of two types of people: those more likely to perceive offensive language and hate speech and those less likely. We create three datasets of simulated annotations which differ in the mix of the annotator types. We then create a fourth dataset, using the PAIR algorithm, to fix the imbalance in the annotators. We fine-tune RoBERTa models on the four datasets and evaluate the effect of annotator composition on model performance (RQ1) and the ability of the PAIR algorithm to improve performance (RQ2).
4.1 Simulating Annotations
We use our previously collected dataset on tweet annotation sensitivity (Kern et al., 2023) [^3] https://huggingface.co/datasets/soda-lmu/tweet-annotation-sensitivity-2, which is a dataset of 3,000 English-language tweets, each with 15 annotations of both offensive language (OL: yes/no) and hate speech (HS: yes/no). We chose this dataset because the high number of annotations of each tweet gives us a diverse set of labels to work with. We randomly select (without replacement) 12 annotations (of both OL and HS) of each tweet in the original dataset. [^4] As shown in Table 1, we can more carefully control the construction of our datasets when the number of annotations per tweet is even. Let $p_{i,OL}$ be the proportion of the 12 annotators who annotated tweet $i$ as OL and $p_{i,HS}$ defined similarly. Figure 2 shows the distribution of these proportions across the 3,000 tweets. The HS annotations are clustered near 0, whereas the OL annotations are more spread out between 0 and $1$.
Figure 2: Distribution of $p_{i,OL}$ and $p_{i,HS}$ in original data
The population contains two types of people (50% each). Type A people are less likely to say a tweet contains OL. Type B people are more likely:
$p_{i,OL}^{A}$ $=\text{max}(p_{i,OL}-\beta,0)$ (4)
$p_{i,OL}^{B}$ $=\text{min}(p_{i,OL}+\beta,1)$ (5)
Here $\beta$ captures the magnitude of the bias. We vary $\beta$ from $[0.05,0.3]$ by $0.05$, corresponding to an increase or decrease in the probability to judge a tweet as OL by five to 30 percentage points. This range is large on the probability scale and covers most reasonable situations. With these six values of $\beta$, we create six vectors of probabilities ($p_{i,OL}^{A},p_{i,OL}^{B}$) for each tweet.
We then create four datasets, each with 3,000 tweets (Table 1), for each value of $\beta$. The Representative Dataset contains OL annotations from six A annotators (drawn from $\text{Bernoulli}(p_{i,OL}^{A})$) and six B annotators (drawn from $\text{Bernoulli}(p_{i,OL}^{B})$). The proportion of A and B annotators in this dataset matches the simulated population we created.
Annotations A B
Dataset per tweet annotations annotations
Representative 12 6 6
Non-representative 1 9 6 3
Non-representative 2 12 9 3
Adjusted 12 6 3 + 3*
* 3 B annotations replicated
Table 1: Four training datasets for each bias value ($\beta$)
We next create two unbalanced datasets. Non-representative 1 randomly deletes three B annotations for each tweet from the Representative Dataset. Non-representative 2 adds three additional A annotations, drawn from $p_{i,OL}^{A}$, to the Non-representative 1 dataset. The Non-representative 2 Dataset is more unbalanced than Non-representative 1, but contains the same number of annotations as the Representative dataset.
4.2 Applying PAIR Algorithm
Finally, we use the PAIR algorithm to create the Adjusted Dataset. Starting with the Non-representative 1 Dataset, we calculate the share of the annotator pool that is in the A and B strata: $S_{A}=\frac{2}{3},S_{B}=\frac{1}{3}$. The population proportions, by construction, are $P_{A}=0.5,P_{B}=0.5$. Applying (1), we get $w_{A,i}=0.75,w_{B,i}=1.5$. We multiply these weights by $K=\frac{4}{3}$ to get $w_{A,i}=1,w_{B,i}=2$. These weights give us $n_{A,i}=0,n_{B,i}=1$, which leads us to replicate all B annotations in the Non-representative 1 Dataset (see Table 1).
The HS probabilities for the A and B annotators are defined in the same way: $p_{i,HS}^{A}=\text{max}(p_{i,HS}-\beta,0)$, $p_{i,HS}^{B}=\text{min}(p_{i,HS}+\beta,1)$. We also construct the four datasets (Representative, Non-representative 1, Non-representative 2, Adjusted) in the same way we did in the OL case.
Figures 3 and 4 show the percentage of instances annotated OL and HS in the four datasets for each value of $\beta$. In both, the percentage of OL/HS annotations in the Adjusted dataset is similar to that in the Representative dataset for all values of $\beta$. The percentage in the two unbalanced datasets is lower, because those datasets overrepresent the A annotators, who are less likely to annotate OL/HS.
HS is rare in our dataset (16.7% of instances were annotated as HS), and our simulation strategy overrepresents A annotators in the two Non-representative datasets, who are less likely to perceive HS (Table 1). For these reasons, as $\beta$ increases, more $p_{i,HS}^{A}$ are 0 while the $p_{i,HS}^{B}$ probabilities increase. This issue leads the proportion of HS annotations in the Representative and Adjusted datasets to increase with $\beta$ in the HS dataset, which have more B annotations than the unadjusted datasets (Figure 4).
Figure 3: Percentage of instances annotated as OL, by dataset and bias ($\beta$)
Figure 4: Percentage of instances annotated as HS, by dataset and bias ($\beta$)
4.3 Model Training and Evaluation
Training and Test Setup.
We train models on each of the eight datasets: four for OL, four for HS. We divide each dataset, at the tweet level, into training (2,000 tweets), development (500), and test (500) sets. Each tweet appears 12 times in the Representative, Non-representative 2, and Adjusted datasets and nine times in the Non-representative 1 set.
Model Selection and Training.
We used RoBERTa base Liu et al. (2019) as our text classifier, training for five epochs on each dataset, with development set optimization. To ensure reliable results, we trained five versions with different random seeds and averaged their performance.
Our implementation of RoBERTa models was based on the libraries pytorch (Paszke et al., 2019) and transformers (Wolf et al., 2020). During training, we used the same hyperparameter settings (Table 2) for the five training conditions to keep these variables consistent for comparison purposes. We trained each model variation with five random seeds $\{10,42,512,1010,3344\}$ and took the average across the models. All experiments were conducted on an NVIDIA® A100 80 GB RAM GPU.
Hyperparameter Value
encoder roberta-base
epochs_trained 55
learning_rate $3e^{-5}$
batch_size $32$
warmup_steps $500$
optimizer AdamW
max_length $128$
Table 2: Hyperparameter settings of RoBERTa models
Performance Metrics.
We evaluate models using calibration and accuracy metrics on the test set. While accuracy metrics directly measure classification performance, calibration metrics provide crucial insights into model reliability by assessing probability estimate quality – particularly important for high-stakes applications requiring trustworthy confidence measures.
For calibration, we report Absolute Calibration Bias (ACB, Equation 6), which measures how well a model’s predicted probabilities align with true annotation frequencies. For each tweet $i$, we compare the model’s predicted probability of offensive language ($\text{preds}_{i,OL}$) against the true proportion of annotators who labeled that tweet as offensive ($p_{i,OL}$).
$\text{ACB}_{OL}=\frac{1}{n}\sum_{i=1}^{n}\left|\text{preds}_{i,OL}-p_{i,OL}\right|$ (6)
$\text{ACB}_{HS}$ is defined accordingly. ACB adapts established calibration metrics by using the annotator agreement proportion as a plug-in estimator for the true probability, avoiding the need for binning (as in ECE, Naeini et al., 2015) while maintaining the intuitive L1 distance interpretation (Roelofs et al., 2022). A low ACB score indicates that the model’s confidence scores accurately reflect the underlying annotation uncertainty in the population.
For accuracy, we report the F1 score.
5 Results
We show results separately for the OL and HS models.
5.1 OL Models
Calibration.
Figure 5 compares the ACB in the test set for models trained on the four datasets. The dark lines show average ACB across the five training runs and the shading shows the standard deviation.
The ACB for the models trained on the Adjusted dataset closely tracks that for the Representative dataset and does not increase with $\beta$. ACB for the models trained on the two unbalanced datasets is greater and grows with $\beta$. These results demonstrate the effectiveness of our adjustment method. Replicating the annotations from the underrepresented annotator type to match population proportions improves model calibration.
Figure 5: ACB scores for OL Models, by dataset and bias ($\beta$)
Figure 6: F1 scores for OL Models, by dataset and bias ($\beta$)
Figure 7: ACB scores for HS Models, by dataset and bias ($\beta$)
Figure 8: F1 scores for HS Models, by dataset and bias ($\beta$)
Figure 9: Percentage of OL instances on difficult tweets ($0.4\leq p_{i,OL}\leq 0.6$) by dataset and bias ($\beta$)
Figure 10: Percentage of HS instances on difficult tweets ($0.4\leq p_{i,HS}\leq 0.6$) by dataset and bias ($\beta$)
Accuracy.
Figure 6 compares the models’ F1 scores. In contrast to Figure 5, we do not see strong differences between the models trained on the different datasets. For all datasets, model performance declines with $\beta$: as the amount of bias in the annotations increases, the models are less able to predict the binary OL label.
Because the F1 metric focuses on binary predictions, it is less sensitive to training biases than calibration metrics like ACB, which more explicitly capture biases through prediction scores. In decision-making, miscalibrated predictions can have harmful consequences when, for example, hateful content remains undetected Van Calster et al. (2019). These findings suggest that calibration metrics provide a clearer view of the impact of annotators on models: binary classification metrics can obscure such effects.
5.2 HS Models
Figure 7 contains the ACB results and Figure 8 the F1 score results for the HS models trained on each dataset. Though the adjusted model roughly tracks the representative models for ACB, there is instability in the results. All models show lower average ACB values than the representative model across a wide range of the bias offset (0.10 - 0.20). The PAIR approach does not improve calibration or accuracy: the adjusted model performs similarly to the Non-representative models. This effect is likely due to the combination of label rarity and our simulation design. With few positive annotations to begin with, the impact of the $\beta$ parameter and the overrepresentation of the A annotators may be overwhelmed by the baseline scarcity of hate speech annotations. Calibration metrics can be less reliable with rare classes Zhong et al. (2021).
5.3 Sensitivity Analysis: Difficult Tweets
Our simulations assumed that all annotator type impacts all tweets the same way (Eq. 5), which is an oversimplification. More likely, annotator characteristics have more impact for ambiguous tweets. For example, prior research in the psychology literature on judgment under uncertainty suggests that people draw more heavily on personal heuristics when interpreting unclear or underspecified information Tversky and Kahneman (1974). For this reason, we repeat model training and recompute metrics for those tweets where $0.4\leq p_{i}\leq 0.6$. Subsetting the tweets in this way also eliminates the floor and ceiling effects in Eq. 5. The filtered datasets contain 267 (OL) and 360 (HS) tweets. The proportions of OL and HS annotations are stable for the Representative and Adjusted sets and decrease for the Non-representative sets as we increase the bias offset (Figures 9 and 10). This mimics the trend for OL in the full dataset (Figure 3).
Figure 11: ACB scores for OL Models, on difficult tweets ($0.4\leq p_{i,OL}\leq 0.6$), by dataset and bias ($\beta$)
Figure 12: F1 scores for OL Models, on difficult tweets ($0.4\leq p_{i,OL}\leq 0.6$), by dataset and bias ($\beta$)
Figures 11, 12, 13, and 14 show results for two metrics (ACB, F1) for filtered OL and HS annotations. In Figure 11, the Representative and Adjusted models have similar ACB and are lower than the Non-representative models. The F1 scores do not show differences between the models. These results are similar to those on the full set of tweets (Figures 5 and 6). In the two HS figures (13, 14), we see signs that the Representative and Adjusted models perform similarly, and better than the two Non-representative models, on both metrics. These results are more promising than those on the full set of tweets (Figures 7 and 8) and support our hypothesis that the rarity of HS annotations contributed to the lack of positive results for the PAIR approach in §5.2. The PAIR algorithm works well with difficult tweets, which is where it is likely most needed.
Figure 13: ACB scores for HS Models, on difficult tweets ($0.4\leq p_{i,HS}\leq 0.6$), by dataset and bias ($\beta$)
Figure 14: F1 scores for HS Models, on difficult tweets ($0.4\leq p_{i,HS}\leq 0.6$), by dataset and bias ($\beta$)
6 Discussion & Recommendations
Our experimental results show the OL prediction models perform less well when trained on data from non-representative annotator pools (RQ1), and simple statistical adjustments can improve model calibration without collecting additional annotations or involving additional annotators (RQ2). While PAIR’s impact was harder to assess for the rare HS class, PAIR did improve calibration of both the OL and HS models when trained on difficult tweets. These findings establish a promising bridge between survey statistics and machine learning – offering a practical approach to make AI systems more representative of and responsive to the populations they serve, particularly for tasks involving subjective human judgments.
We recommend the following four steps to reduce bias due to non-representative annotator pools:
  1. 1)
    Use social science research to identify the annotator characteristics that influence the propensity to engage in annotation and the annotations provided Eckman et al. (2024).
  2. 2)
    Collect these characteristics from annotators and gather corresponding population-level data from national censuses or high-quality surveys. [^5] Collection and release of annotator characteristics or weights derived from them may raise confidentiality concerns. The survey literature offers advice for sharing sensitive data (see Karr, 2016, for a review). Collecting annotator characteristics may also require involvement of Institutional Review Boards or other participant protection organizations Kaushik et al. (2024).
  3. 3)
    Calculate weights that match the annotators to the population on those characteristics (Bethlehem et al., 2011; Valliant et al., 2013).
  4. 4)
    Use these weights in model training. Our simple replication approach showed promise, and future work should test more sophisticated weighting approaches.
7 Limitations
Our study explores bias-aware data simulation and evaluation in a controlled setting, which necessarily involves simplifying assumptions and methodological constraints. We outline key areas where future work could broaden the applicability and robustness of our findings.
Stylized Biases and Simulated Data.
Our simulation makes strong assumptions about annotator behavior: there are only two types of annotators, and, within each type, annotators behave similarly. Real-world annotator biases may be more nuanced or context-dependent. The simulated annotators might not be representative of a stable opinion group Mokhberian et al. (2024); Vitsakis et al. (2024). Future work could incorporate more realistic biases and refine the proposed simulations and statistical techniques.
Sampling Variability.
We have created only one version of the four datasets for each annotation type and value of $\beta$, each of which contains random draws from the Bernoulli distribution. A more traditional statistical approach would create multiple versions of the datasets and train models on each one, to average over the sampling variability. Though limited by computational constraints in this work, future work could take on a more expansive simulation. As discussed, we used five seeds in model training.
Need for Population Benchmarks and Annotator Characteristics.
PAIR requires high quality benchmark information about the relevant population. These benchmarks might come from national statistical offices or national surveys. Annotators must provide accurate data on the same characteristics available in the benchmark data. Unfortunately, annotators sometimes do not provide accurate information Chandler and Paolacci (2017); Huang et al. (2023). In addition, theory demonstrates that bias will be reduced only when the characteristics used in weighting correlate with the annotations (Eckman et al., 2024). In our simulation, differences in annotations were driven solely by group membership ($A$, $B$). In the real world, it is challenging to know what characteristics impact annotation behavior for a given task and to find good benchmarks for those characteristics.
Generalization Beyond Task Types.
The study focuses only on binary classification tasks. Many real-world annotation tasks involve multiple classes or labels, which may show different bias patterns. Additional research is needed to extend these methods to more complex classification scenarios.
Evaluation Metrics.
While we measured calibration and accuracy, we did not examine other important metrics such as fairness across subgroups or robustness to adversarial examples. Future work on training data adjustment should assess a broader range of performance measures.
8 Ethical Considerations
In this simulation study, we experiment on a publicly available dataset collected in our previous study Kern et al. (2023), which contains offensive and hateful tweets. We do not support the views expressed in these tweets. The simulation study itself does not collect any new data or raise any ethical considerations.
Acknowledgments
We acknowledge use of the Claude model to edit the text of the paper and to assist in coding. We thank the members of SODA Lab and MaiNLP labs from LMU Munich, and the members of the Social Data Science Group from University of Mannheim for their constructive feedback. This research is partially supported by RTI International, MCML, and BERD@NFDI. BP is supported by ERC Consolidator Grant DIALECT (101043235).